OpenStudy (anonymous):

what is the equation of a line passing through -3,7 and having a slope of -1/5

pooja195 (pooja195):

in slope intercept form?

OpenStudy (anonymous):

yes

pooja195 (pooja195):

\[\huge~\rm~y-y_1=m(x-x_1)\] m=slope x1=x coordinate y1=y coordinate

pooja195 (pooja195):

-3=x1 7=y1 -1/5=m plug them in

OpenStudy (anonymous):

options y=-5x+22 y=5x+22 y=-1/5x+32/5 y=-1/5x+32/5 y=-1/5x+22

pooja195 (pooja195):

Ye we'll get to that

pooja195 (pooja195):

\[\huge~\rm~y-7=-1/5(x-(-3))\] -3=x1 7=y1 -1/5=m \[\huge~\rm~y-7=-1/5(x+3)\] can you isolate y from here?

pooja195 (pooja195):

start by distributing −1/5(x+3)

OpenStudy (anonymous):

I don't know how to do that?

OpenStudy (anonymous):

-1/5x+ -1/5(3)

OpenStudy (anonymous):

-1/5x+-0.6

OpenStudy (anonymous):

help

OpenStudy (anonymous):

what is the question @Kaelyn78

OpenStudy (anonymous):

nevermind I worked through it, thank you anyway